CONFORMATIONAL PROPERTIES, SPECTROSCOPY AND STRUCTURE OF ISATIN-(WATER)_{n=1−3} CLUSTERS

MILIND K. SINGH, D. M. UPADHYA and \underline{VIPIN B. SINGH}, Department of Physics, U P Autonomous College, Varanasi-221002, India.

The structure, stability and vibrational characteristics of Isatin-(Water)_{n} clusters with n=1=3 have been investigated using second order Moller-Plesset (MP2) perturbation tehory and Density Functional Theory (with B3LYP) methods employing the basis set 6-31+G(d). The vertical excitation energies for these complexes have been also computed using the time-dependent density functional theory. The three stable conformational isomers, each for Isatin-(Water)_{1} and Isatin-(Water)_{2} clusters were obtained. It is shown that in the most stable isomer of Isatin-(Water)_{1} cluster hydrogen bond between amide hydrogen and oxygen of water is found stronger as compared to the H-bond in Indole-(Water)_{1} cluster. For a particular position of complexation of water, between the carbonyl oxygen’s, results an unusual increase in the dipole moment due to an electronic charge displacement from the N atom to the C atom of the neighboring carbonyl bond. This causes a large separation between the effective charges forming the dipole. The complexes involving this position of water are expected to show a small charge transfer character. The experimentally observed electronic absorption peaks are reasonably reproduced by the TD-DFT calculations and it is found that the longest wavelength absorption peak of isatin at 406 nm is significantly red shifted after addition of a water molecule.